How much should parents invest today at 6.24% compounded monthly, in order to have 65,000 for their child's educations in 18 years from now? I would appreciate a step by step walkthrough! Thanks!
Answer by mananth(16946)

Principal P = x
Amount= 65000
years=n 18
compounded 12 times a year
Rate = 6.24 0.520
Amount = P*((n+r)/n)^n

65000 = x*(1+r/n)^n


65000 = P(1+0.0624/12 )^12*18

65000 = P( 1.00520 )^ 216

65000 /( 1.01 )^ 216 = P

P= 21202.04